Reducible to homogeneous differential equations or by simple substitution.

1. (2๐‘ฅ โˆ’ 3๐‘ฆ โˆ’ 4)๐‘‘๐‘ฅ โˆ’ (3๐‘ฅ โˆ’ 4๐‘ฆ โˆ’ 2)๐‘‘๐‘ฆ = 0

2. (2๐‘ฅ โˆ’ ๐‘ฆ โˆ’ 3)๐‘‘๐‘ฅ โˆ’ (๐‘ฅ + 4๐‘ฆ + 3)๐‘‘๐‘ฆ = 0

3. (๐‘ฅ โˆ’ ๐‘ฆ โˆ’ 6)๐‘‘๐‘ฆ = (๐‘ฅ โˆ’ ๐‘ฆ + 2)๐‘‘๐‘ฅ

4. (๐‘ฅ โˆ’ 2๐‘ฆ + 4)๐‘‘๐‘ฅ + (2๐‘ฅ โˆ’ 4๐‘ฆ โˆ’ 1)๐‘‘๐‘ฆ = 0

5. (๐‘ฅ + 4๐‘ฆ + 3)๐‘‘๐‘ฅ = (2๐‘ฅ โˆ’ ๐‘ฆ โˆ’ 3)๐‘‘๐‘ฆ


B. Exact differential equations (include checking for exactness).

1. (๐‘ค3 + ๐‘ค๐‘ง 2 โˆ’ ๐‘ง)๐‘‘๐‘ค + (๐‘ง 2 + ๐‘ค2 ๐‘ง โˆ’ ๐‘ค)๐‘‘๐‘ง = 0

2. (cos 2๐‘ฆ โˆ’ 3๐‘ฅ 2๐‘ฆ 2 )๐‘‘๐‘ฅ + (cos 2๐‘ฆ โˆ’ 2๐‘ฅ sin 2๐‘ฆ โˆ’ 2๐‘ฅ 3๐‘ฆ)๐‘‘๐‘ฆ = 0


1
Expert's answer
2021-12-07T09:55:59-0500

A.

1.

"x\to X+h, y\to Y+k"

"2h-3k-4=0"

"3h-4k-2=0"

"h=k-2"

"2k-4-3k-4=0"

"k=-8,h=-10"

"x=X-10,y=Y-8"


"\frac{dY}{dX}=\frac{2X-3Y}{3X-4Y}"


"Y=tX,Y'=t'X+t"


"t'X+t=\frac{2X-3tX}{3X-4tX}"


"t'X=\frac{2-3t-3t+4t^2}{3-4t}"


"\frac{3-4t}{4t^2-6t+2}dt=\frac{dX}{X}"


"-\frac{ln(2t^2-3t+1)}{2}=lnX+lnc_1"


"\frac{1}{2t^2-3t+1}=c_2X^2"


"\frac{1}{2(Y/X)^2-3(Y/X)+1}=c_2X^2"


"2(\frac{y+8}{x+10})^2-3\frac{y+8}{x+10}+1=\frac{c}{(x+10)^2}"


2.

"x\to X+h, y\to Y+k"

"2h-k-3=0"

"h+4k+3=0"

"h=5k+6"

"9k+9=0"

"k=-1,h=1"

"x=X+1,y=Y-1"


"\frac{dY}{dX}=\frac{2X-Y}{X+4Y}"


"Y=tX,Y'=t'X+t"


"t'X+t=\frac{2X-tX}{X+4tX}=\frac{2-t}{1+4t}"


"t'X=\frac{-4t^2-2t+2}{1+4t}"


"\frac{1+4t}{-4t^2-2t+2}dt=\frac{dX}{X}"


"-\frac{ln(2t^2+t-1)}{2}=lnX+lnc_1"


"\frac{1}{2t^2+t-1}=c_2X^2"


"\frac{1}{2(Y/X)^2+(Y/X)-1}=c_2X^2"


"2(\frac{y+1}{x-1})^2+\frac{y+1}{x-1}-1=\frac{c}{(x-1)^2}"


3.

"x-y=v"


"1-\frac{dv}{dx}=\frac{v+2}{v-6}"


"-\frac{8}{v-6}=\frac{dv}{dx}"


"-8x=\frac{(v-6)^2}{2}+c"


"-8x=\frac{(x-y-6)^2}{2}+c"


4.

"x-2y=v"


"\frac{1}{2}(1-\frac{dv}{dx})=-\frac{v+4}{2v-1}"


"1+\frac{2v+8}{v-6}=\frac{dv}{dx}"


"\frac{3v+2}{v-6}=\frac{dv}{dx}"


"dx=\frac{v-6}{3v+2}dv"


"x=\frac{3v-20ln(3v+2)}{9}+c"


"x=\frac{3(x-2y)-20ln(3(x-2y)+2)}{9}+c"


5.

"x\to X+h, y\to Y+k"

"2h-k-3=0"

"h+4k+3=0"


"h=5k+6"

"9k+9=0"

"k=-1,h=1"


"x=X+1,y=Y-1"


"\frac{dY}{dX}=\frac{X+4Y}{2X-Y}"


"Y=tX,Y'=t'X+t"


"t'X+t=\frac{X+4tX}{2X-tX}=\frac{1+4t}{2-t}"


"t'X=\frac{-t^2+2t+1}{1+4t}"


"\frac{1+4t}{-t^2+2t+1}dt=\frac{dX}{X}"


"\frac{5}{2\sqrt 2}ln(\frac{t+\sqrt 2-1}{t-\sqrt 2-1})-2ln(t^2-2t-1)=lnX+lnc"


"\frac{5}{2\sqrt 2}ln(\frac{Y/X+\sqrt 2-1}{Y/X-\sqrt 2-1})-2ln((Y/X)^2-2Y/X-1)=ln(cX)"


"\frac{5}{2\sqrt 2}ln(\frac{(y+1)/(x-1)+\sqrt 2-1}{(y+1)/(x-1)-\sqrt 2-1})-2ln(((y+1)/(x-1))^2-2(y+1)/(x-1)-1)="


"=ln(c(x-1))"


B.

1.

"(\ud835\udc64^3 + \ud835\udc64\ud835\udc67 ^2 \u2212 \ud835\udc67)_z=2zw-1"

"(\ud835\udc67 ^2 + \ud835\udc64^2 \ud835\udc67 \u2212 \ud835\udc64)_w=2zw-1"


"F=\int (𝑤^3 + 𝑤𝑧 ^2 − 𝑧)dw=w^4/4+w^2z^2/2-zw+g(z)"


"F_z=w^2z-w+g'(z)=\ud835\udc67 ^2 + \ud835\udc64^2 \ud835\udc67 \u2212 \ud835\udc64"

"g'(z)=\ud835\udc67 ^2"

"g(z)=\int𝑧 ^2dz=z^3/3+c"


"F=w^4\/4+w^2z^2\/2-zw+z^3\/3+c"

"w^4\/4+w^2z^2\/2-zw+z^3\/3+c=0"


2.

"(cos 2\ud835\udc66 \u2212 3\ud835\udc65^ 2\ud835\udc66 ^2 )_y=-2sin2y-6x^2y"

"(cos 2\ud835\udc66 \u2212 2\ud835\udc65 sin 2\ud835\udc66 \u2212 2\ud835\udc65 ^3\ud835\udc66)_x=-2sin2y-6x^2y"


"F=\int (cos 2𝑦 − 3𝑥^ 2𝑦 ^2 )dx=xcos2y-x^3y^2+g(y)"


"F_y=-2xsin2y-2yx^3+g'(y)=cos 2\ud835\udc66 \u2212 2\ud835\udc65 sin 2\ud835\udc66 \u2212 2\ud835\udc65 ^3\ud835\udc66"

"g'(y)=cos 2\ud835\udc66"


"g(y)=\intop cos2ydy=sin2y/2+c"


"F=xcos2y-x^3y^2+sin2y\/2+c"

"xcos2y-x^3y^2+sin2y\/2+c=0"